Let the production function of a firm is given as

𝑞=(𝑥0.5 +𝑦0.5)2

Where 𝑥 and 𝑦 are inputs and 𝑤𝑥 is the price of input 𝑥 and 𝑤𝑦 is the price

of input 𝑦.

a) Assume the firm has a limited budget to spend on buying input. Find

the cost-conditional input demand function for each input.

b) Find the cost function of the firm.


1
Expert's answer
2021-10-12T09:56:50-0400

a .)

Demand condition for the factors of production is achieved at the point where the production is at the optimal level for a given price ratio of the factors . 

Optimal Production condition for a firm is where the MRTS is equal to the factor price ratio .

& MRTS = Marginal Rate of technical Substitution "=\frac{MU(x) }{MU(y)}"  

In our example , two inputs : x & y 

Price of x = Wx 

Price of y = Wy 

Hence the Costs will be in the form :

"Wx\times x + Wy\times y = C"  

Factor price ratio "= \frac{Wx}{Wy}"

 


Production function : "q = 2 (x^{0.5} + y^{0.5} )"

"MP(x) = \frac{dq}{dx} = x\\

MP(y) =\frac{ dq}{dy} = y\\ 

MRTS = \frac{x}{y}"

Optimal Condition : MRTS = factor Price Ratio 

"\frac{x}{y} = \frac{Wx}{Wy }\\

x = y×\frac{Wx}

{Wy}

-----(i)"

Putting this value of (x) into the Cost condition faced by producer :

"Wx\times x + Wy\times y = C \\

Wx (y\times \frac{Wx}{

Wy}



 + Wy\times y = C \\

y* = \frac{C

}{\frac{W^2}{Wy}+Wy}\\


y*=\frac{Wy×C}{Wx^2 + Wy^2}"    (Demand function for input y )

Putting this value of y* into the optimal equation (i) 

"x* = y\times ×\frac{Wx}{Wy}\\

x*=\frac{Wx×C}{Wx^2 + Wy^2}" (Demand function for input y ) 


b .)

Cost function for a given technology refers to the minimum expenditure which would be made to achieve certain production level . 

Hence , cost function could be depicted as the optimal cost that would be incurred . So , we put demand functions into the cost condition for firm :

Cost Function "= Wx(x*) + Wy(y*)"

"= Wx (\frac{Wx×C}{Wx^2 + Wy^2}) + Wy(\frac{Wy×C}{Wx^2 + Wy^2}\\
= \frac{C}{Wx^2+Wy^2}(Wx^2 + Wy^2)"

Cost function = C = Wx(x) + Wy(y)